OpenStudy (dtan5457):

What did I do wrong while trying to solve this complex fraction? (I'll show all my work)

OpenStudy (alexandervonhumboldt2):

k

OpenStudy (alexandervonhumboldt2):

i will help him @Nnesha

OpenStudy (dtan5457):

\[\frac{ \frac{ 1 }{ z-7 }+\frac{ 7 }{ z } }{ \frac{ z }{ z^2-7z }-1 }\]

OpenStudy (dtan5457):

I multiplied the numerator by x(x-7)

OpenStudy (dtan5457):

and got \[\frac{ 8z-49 }{z^2-7z } \]

OpenStudy (dtan5457):

Then for the bottom part i multiplied the -1 by the same thing : x(x-7)

OpenStudy (dtan5457):

I got\[\frac{ z }{ z^2-7z }-\frac{ -z^2+7z }{ z^2-7z }=\frac{ -z^2-6z }{ z^2-7z }\]

Nnesha (nnesha):

hey wait do you mean multiply by z(z-7)

OpenStudy (dtan5457):

I put these on top of each other and flipped the 2nd term to get \[\frac{ 8z-49 }{ z^2-7z }\times \frac{ z^2-7z }{ -z^2-6z }\]

OpenStudy (dtan5457):

and yes

OpenStudy (dtan5457):

now i'm stuck with \[\frac{ 8z-49 }{ -z^2-6z }\]

OpenStudy (dtan5457):

which isn't the right answer

OpenStudy (dtan5457):

based on my work, what did i do wrong?

OpenStudy (dtan5457):

someone please enlighten me ive been stuck on this for a few hours

OpenStudy (misty1212):

\[\frac{ \frac{ 1 }{ z-7 }+\frac{ 7 }{ z } }{ \frac{ z }{ z^2-7z }-1 }\times \frac{z(z-7)}{z(z-7)}\] \[=\frac{z+7(z-7)}{z-z(z-7)}\]

Parth (parthkohli):

\[\frac{ \frac{ 1 }{ z-7 }+\frac{ 7 }{ z } }{ \frac{ z }{ z^2-7z }-1 }\]\[= \dfrac{\frac{z + 7 (z - 7)}{z(z-7)}}{\frac{z - z(z - 7)}{z(z - 7)}} \]\[= \dfrac{z + 7(z - 7)}{z - z(z -7)}\]\[= \dfrac{z + 7z - 49}{z - z^2 + 7z}\]\[= \dfrac{8z - 49}{ 8z-z^2}\]

OpenStudy (misty1212):

remove the parentheses in the numerator and the denominator get \[\frac{z+7z-49}{z-z^2+7}\] cleans up to \[\frac{8z-49}{-z^2+z+7}\]

OpenStudy (dtan5457):

i multiplied the numerator part only the 2nd term by z(z-7) and the first term by z

OpenStudy (dtan5457):

so they have the same denominator?

OpenStudy (misty1212):

oh i made a mistake @ParthKohli has it better then me

OpenStudy (misty1212):

\[\frac{8z-49}{-z^2+z+7}\] is wrong it is \[\frac{8z-49}{-z^2+z+7z}\]

OpenStudy (dtan5457):

well i got the numerator correct

OpenStudy (dtan5457):

but the bottom part i have -z^2-6z

OpenStudy (misty1212):

before the cleaning up the denominator is what @ParthKohli wrote \[z-z(z-7)\]

OpenStudy (dtan5457):

I don't know is this correct for the denominator? \[\frac{ z }{ z^2-7z }-1=\frac{ -z^2+7z }{ z^2-7z }\]

OpenStudy (dtan5457):

I put -1 as -1/1 then multiplied that with z^2-7z

OpenStudy (dtan5457):

so now \[\frac{ z }{ z^2-7z }-\frac{ -z^2+7z }{ z^2-7z }=\frac{ -z^2-6z }{ z^2-7z }??\]

OpenStudy (dtan5457):

@Jhannybean

OpenStudy (dtan5457):

@saifoo.khan

OpenStudy (dtan5457):

@DanJS

OpenStudy (dtan5457):

@AlexandervonHumboldt2

OpenStudy (paxpolaris):

\[\frac{ z }{ z^2-7z }-\frac{ -z^2+7z }{ z^2-7z }=\frac{ -z^2-6z }{ z^2-7z }\] you are doing .... \[\frac{ z }{ z^2-7z }-(-1)\] instead of doing: \[\frac{ z }{ z^2-7z }-1\] you are taking the negative twice which makes it wrong

OpenStudy (dtan5457):

if i multiply -1 by z^2-7z won't that change the signs?

OpenStudy (paxpolaris):

\[\frac{ z }{ z^2-7z }-1\\=\frac{ z }{ z^2-7z }+(-1)\\=\frac{ z }{ z^2-7z }+{(-1)(z^2-7z)\over z^2-7z}\\=\frac{ z }{ z^2-7z }\color{red}+{\color{red}-z^2\color{red}+7z \over z^2-7z}\] \[=\large{\color{red}-z^2\color{red}{+8}z \over z^2-7z}\]

OpenStudy (dtan5457):

o.

OpenStudy (dtan5457):

so if i can take the new numerator and denominator i get \[\frac{ 8z-49 }{ -z^2+8z }\]

OpenStudy (paxpolaris):

\[\checkmark\]

OpenStudy (dtan5457):

on my regents book the numerator is 7z-47..

OpenStudy (dtan5457):

the denominator is correct

OpenStudy (dtan5457):

They can get that if they subtracted on the numerator but I don't see that being possible?

OpenStudy (paxpolaris):

maybe you are looking at a different question ...to get numerator 7z-47 you have to have \[{2\over z(z-7)} + \frac7z\] ....for the numerator

OpenStudy (dtan5457):

There's is definitely a 1 instead of a 2...

OpenStudy (dtan5457):

Book error???

OpenStudy (paxpolaris):

i guess..

OpenStudy (dtan5457):

that's pretty crazy....most absurd question..seriously.

OpenStudy (dtan5457):

Thanks for clearing up my mistake on the denominator though.

OpenStudy (paxpolaris):

np